If f(x) = 5x - 1, find the following.
f(a-2)

Answer by uknouluvme89(30)

If f(x) = 5x - 1, find the following.
f(a-2) = 5(a-2) -1
f(a-2) = 5a-10-1
f(a-2) = 5a-11